# Reconstruction-Free Anomaly Detection with Diffusion Models

### v2

@@ Conclusion
We utilize three publicly available anomaly detection (AD) benchmarks for evaluation: MVTecAD (Bergmann et al. 2019), VisA (Zou et al. 2022), MPDD (Jezek et al. 2021). Table 8 summarizes the key statistics of the datasets used in our experiments, including the number of classes, domain, image resolution, and sample counts for training and testing. All of these datasets follow an unsupervised AD setting. In other words, the training set only consists of defect-free normal images, and the test set consists of both normal and anomalous images.
Appendix B B. Implementation Details
Data Preparation
@@
Model Setup
We adopt the original DiT architecture (Peebles and Xie 2023), but remove all class‐conditioning so that our method does not depend on class labels. This design allows DiT to be applied in scenarios where ground‐truth annotations are unavailable or impractical (e.g., inspection of diverse products on a manufacturing line). We set the number of transformer blocks and embedding dimensions to 16 and 2048 for DiT-gigant, and to 8 and 1024 for DiT-base, respectively. For the UNet‐based diffusion models evaluated in Table 7, we follow the architectural specifications of the original DDPM implementation (Ho, Jain, and Abbeel 2020), consistent with prior work (Beizaee et al. 2025; He et al. 2024). For the MLP‐based diffusion models, we stack multiple MLP blocks and apply AdaLN-Zero conditioning (Peebles and Xie 2023), as in (Li et al. 2025). We summarize all major hyperparameters for diffusion modeling in Table 9.
Table 8: Dataset statistics used in our experiments. ‘N’ and ‘A’ denote the numbers of normal and anomalous samples, respectively.
MVTecAD VisA MPDD
Classes 15 12 6
Domain Industrial Industrial Metal
Resolution $\times1024$ $\times256$ $\times1024$
Train (N) 2,629 8,659 888
Test (N) 467 962 176
Test (A) 1,258 1,200 282
@@
  • Area Under the Per-Region-Overlap Curve (AU-PRO) calculates the IoU between prediction and anomalous region, which is more robust to the small anomalies (Bergmann et al. 2019).
Experimental Environment
Our implementations are mostly based on Pytorch-2.7.1 and Python-3.10. We show the machine specification for the training and evaluation in our experiments in Table 10.
@@
Appendix C C. Additional Quantitative Results on Industrial Anomaly Detection
To further investigate the effectiveness of our proposed framework, we evaluate in Table 11 the detection performance on VisA (Zou et al. 2022) in a single-class setting. In addition to the state-of-the-art performance on the multi-class AD setting explained in our manuscript (Table 3), our InvAD also shows competitive single-class detection accuracy with a remarkable efficiency gain.
Table 11: Quantitative results per category on VisA under the single-class setting, where we report (image-level AU-ROC, pixel-level AU-ROC) and FPS. The best results per category are highlighted in bold.
Category GLAD TransFusion DAD DDAD InvAD (Ours)
candle (99.9, 94.8) (98.3, N/A) (98.7, N/A) (99.9, 98.7) (98.0, 99.4)
capsules (99.1, 99.6) (99.6, N/A) (97.9, N/A) (100, 99.5) (92.1, 99.7)
cashew (98.4, 97.0) (93.7, N/A) (96.5, N/A) (94.5, 97.4) (95.6, 98.9)
chewinggum (99.6, 99.1) (99.6, N/A) (99.9, N/A) (98.1, 96.5) (99.7, 98.4)
fryum (99.4, 96.9) (98.3, N/A) (98.3, N/A) (99.0, 96.9) (97.5, 96.9)
macaroni1 (99.9, 99.8) (99.4, N/A) (99.5, N/A) (99.2, 98.7) (97.8, 99.7)
macaroni2 (98.9, 99.8) (96.5, N/A) (99.0, N/A) (99.2, 98.2) (86.9, 99.4)
pcb1 (99.6, 99.6) (98.9, N/A) (99.2, N/A) (100, 93.4) (99.5, 99.8)
pcb2 (100, 98.6) (99.7, N/A) (99.1, N/A) (99.7, 97.4) (99.5, 99.1)
pcb3 (99.9, 98.9) (99.2, N/A) (98.6, N/A) (97.2, 96.3) (98.6, 99.2)
pcb4 (99.9, 99.5) (99.6, N/A) (98.9, N/A) (100, 98.5) (99.8, 98.3)
pipe fryum (98.9, 99.4) (99.6, N/A) (99.8, N/A) (100, 99.5) (99.1, 99.2)
Avg. (99.5, 98.6) (98.5, 88.8) (98.8, 98.9) (98.9, 97.6) (97.0, 99.0)
FPS 0.2 1.6 9.1 0.8 88.1
Appendix D D. Generalizability on Other Domains
While our InvAD achieves state-of-the-art performance on industrial image AD benchmarks (e.g., MVTecAD, VisA, MPDD), it is still unclear whether it generalizes to other AD tasks like medical imaging. To address this, we evaluate InvAD on a representative medical imaging benchmark, BMAD (Bao et al. 2024), which is a comprehensive AD benchmark including six reorganized datasets supporting medical AD domains, such as brain MRI, liver CT, and retinal OCT. The BMAD consists of 52,742 normal training images, 22,632 anomalous test images, and 4,540 normal test images. We report image-level AU-ROC (det.), and pixel-level AU-ROC (loc.), if the ground-truth anomaly maps are provided. For the comparison with other AD methods, we directly report the results in BMAD (Bao et al. 2024).
We compare quantitative results per dataset on BMAD under the single-class setting in LABEL:tab:sup_bmad. In summary, our proposed InvAD achieves state-of-the-art performance on the BMAD benchmark with huge improvements in inference efficiency. These results further support the effectiveness of our proposed framework, not only within the domain of industrial inspection but also in other fields, such as medical diagnosis.
Table 12: Quantitative results on BMAD under the single-class setting, where we report image-level AU-ROC (det.), pixel-level AU-ROC (loc.), and FPS. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics. FPS measures the image-level detection efficiency on the RESC dataset.
Methods BraTS2021 BTCV + LiTS RESC OCT2017 RSNA Camelyon16 mAD FPS
det. loc. det. loc. loc. loc. det. det. det.
PaDiM [ICPR’20] 79.0 94.4 50.8 90.9 75.9 91.4 91.8 77.5 67.3 79.9 20.4
CFlow [CVPR’20] 74.8 93.8 50.8 92.4 75.0 93.8 85.4 71.5 55.7 77.0 15.1
RD4AD [CVPR’22] 89.5 96.5 60.4 96.0 87.8 96.2 97.3 67.6 66.8 84.2 22.6
PatchCore [CVPR’22] 91.7 97.0 60.3 96.4 91.6 96.5 98.6 76.1 69.3 86.4 20.1
SimpleNet [CVPR’23] 82.5 94.8 72.3 97.5 76.2 77.1 94.7 69.1 62.4 80.7 10.8
InvAD (Ours) 90.3 97.8 60.6 97.8 93.6 96.5 99.2 82.7 67.7 87.4 64.3
Appendix E E. Additional Background of Diffusion Models
Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPM) (Ho, Jain, and Abbeel 2020) are a class of latent generative models which involve progressively collapsed $T$ random variables $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$. Let us assume we have i.i.d. data samples $\mathcal{X}=\{\mathbf{x}_{0}\mid\mathbf{x}_{0}\sim q_{0}(\mathbf{x})\}$ where $q_{0}(\mathbf{x})$ is an unknown data distribution. DDPM aims to approximate this data distribution by tracing the data collapsing process (forward process) in reverse order. In general, the forward process $q\bigl{(}\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\bigr{)}$ is represented as a Markov process with the Gaussian transition kernel:
@@
Denoising Diffusion Implicit Models
The forward process in Equation 9 is a Markov process where the current state $\mathbf{x}_{t}$ solely depends on the previous state $\mathbf{x}_{t-1}$. Denoising diffusion implicit models (DDIM) (Song, Meng, and Ermon 2021) is a non-Markov diffusion model where each forward process is conditioned on both $\mathbf{x}_{0}$ and $\mathbf{x}_{t-1}$. Specifically, the forward posterior is defined as:
@@
where $\bm{f}_{\theta}(\mathbf{x}_{t},t)=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$. For more detailed proof and derivation of DDIM, see (Song, Meng, and Ermon 2021).
Derivation of DDIM Inversion
@@
Appendix F F. Ablation on Scoring Scheme
As explained in the paragraph “Anomaly Score Calculation” of our manuscript, our anomaly scoring scheme considers both negative-log-likelihood (NLL) in the noise prior and the spatial absolute difference of Euclidean norm (Abs) to alleviate the reverse-scoring problem (Nalisnick et al. 2019; Osada et al. 2023). To study the advantage of this strategy, we compare the performance in Table 13 with the conventional NLL-based scoring on MVTecAD. It can be seen that our scoring scheme (i.e., NLL + Abs) contributes to robustness to the diffusion step setting $S$, while the NLL or Abs solely do not achieve this. This demonstrates that our scoring scheme plays a key role in the proposed InvAD paradigm to realize reconstruction-free AD, where $S$ can be further set to be very small for efficiency.
Table 13: Effect of the scoring scheme, under different total diffusion steps $S$ in multi-class MVTecAD with image-lavel AU-ROC.
Total Diffusion Steps $S$ 3 5 10 50 100 1000
NLL 96.1 95.2 93.0 89.7 89.4 89.1
Abs 98.4 98.1 95.8 83.2 82.2 82.0
NLL + Abs (Ours) 99.0 98.9 98.4 96.0 95.7 95.4
Appendix G G. Additional Visualizations
We exhibit more visualization against qualitative AD results on MVTecAD and VisA, in Figure 3 and Figure 4.
Figure 3: Additional visualization of anomaly localization on MVTecAD.
Figure 4: Additional visualization of anomaly localization on VisA.

### v3

@@ 6 Dataset Details
We utilize four publicly available anomaly detection (AD) benchmarks for evaluation: MVTecAD [MVTecAD], VisA [ViSA], MPDD [MPDD], and BMAD [bmad]. Table 10 summarizes the key statistics of the datasets used in our experiments, including the number of classes, domain, image resolution, and sample counts for training and testing. All of these datasets follow an unsupervised AD setting. In other words, the training set only consists of defect-free normal images, and the test set consists of both normal and anomalous images.
7 Implementation Details
7.1 Data Preparation
@@
7.2 Model Setup
We adopt the original DiT architecture [DiT] as the backbone of our diffusion modeling, but remove all class‐conditioning so that our method does not depend on class labels. This design allows DiT to be applied in scenarios where ground‐truth annotations are unavailable or impractical (e.g., inspection of diverse products on a manufacturing line). We set the number of transformer blocks and embedding dimensions to 16 and 2048 for DiT-gigant, and to 8 and 1024 for DiT-base, respectively. For the UNet‐based diffusion models evaluated in Table 8 (in the main paper), we follow the architectural specifications of the original DDPM implementation [DDPM], consistent with prior work [decodiff, DiAD]. For the MLP‐based diffusion models, we stack multiple MLP blocks and apply AdaLN-Zero conditioning [DiT], as in [mar]. We summarize all major hyperparameters for diffusion modeling in Table 11.
7.3 Backbone Setup
For fair comparisons with previous works [omiad, HVQ-Trans], we employ pre-trained EfficientNet-B4 [EfficientNet] as our feature extractor. To capture anomalies at different scales, we first extract feature maps from the 1st to the 4th blocks from EfficientNet-B4, and then apply bilinear interpolation followed by channel-wise concatenation, creating $\mathbf{z}\in\mathbb{R}^{272\times 16\times 16}$, as in [omiad, HVQ-Trans].
We evaluate in Table 7 in the main paper the robustness of our method to the different feature space encoding modules: ViT-B [ViT] and DINO-base [dino]. Both modules employ the vision transformer-based architecture and use ImageNet-1k as the pre-training dataset. The only difference comes from the pre-training scheme (i.e., supervised for ViT-B [ViT], and self-distillation for DINO-base [dino]). For both models, we extract patch tokens of the last block and reorder them to construct feature maps, creating $\mathbf{z}\in\mathbb{R}^{768\times 16\times 16}$.
Table 10: Dataset statistics used in our experiments. ‘Res.’ represents image resolution, where $H=W$. ‘N’ and ‘A’ denote the numbers of normal and anomalous samples, respectively.
MVTecAD VisA MPDD BMAD
Classes 15 12 6 6
Domain Industrial Industrial Metal Medical
Res. 1024 256 1024 1024
Train (N) 2,629 8,659 888 52742
Test (N) 467 962 176 4540
Test (A) 1,258 1,200 282 22632
@@ 7.4 Evaluation Pipeline
  • Area Under the Per-Region-Overlap Curve (AU-PRO) calculates the IoU between prediction and anomalous region, which is more robust to the small anomalies [MVTecAD].
7.5 Experimental Environment
Our implementations are based on Pytorch-2.7.1 [torch] and Python-3.10. We show the machine specification for the training and evaluation in our experiments in Table 12.
@@ 7.5 Experimental Environment
It is important to understand that we do not train any extra conditioning modules nor perform reconstruction, just train unconditional diffusion models in the standard way. This leads to significant training time reduction (e.g., for MVTecAD training time, InvAD: 2.5h on 8×H200 @28 GB/GPU— 2 × faster than DiAD [DiAD]’s 5h). The use of 8× H200 is optional, not required. Specifically, training can be performed on GPUs with more than 32 GB of memory for our largest models.
8 Additional Background of Diffusion Models
8.1 Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPM) [DDPM] are a class of latent generative models which involve progressively collapsed $T$ random variables $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$. Let us assume we have i.i.d. data samples $\mathcal{X}=\{\mathbf{x}_{0}\mid\mathbf{x}_{0}\sim q_{0}(\mathbf{x})\}$ where $q_{0}(\mathbf{x})$ is an unknown data distribution. DDPM aims to approximate this data distribution by tracing the data collapsing process (forward process) in reverse order. In general, the forward process $q\bigl(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\bigr)$ is represented as a Markov process with the Gaussian transition kernel:
@@ 8.1 Denoising Diffusion Probabilistic Models
(a) Histogram of normal and anomalous samples with NLL scoring.
(b) Histogram of normal and anomalous samples with our NLL+Abs scoring.
Figure 4: Comparison of the histogram of normal and anomalous samples, with the conventional NLL scoring (a) and our proposed NLL+Abs scoring (b), on the test set of hazelnut in MVTecAD.
8.2 Denoising Diffusion Implicit Models
The forward process in Equation 9 is a Markov process where the current state $\mathbf{x}_{t}$ solely depends on the previous state $\mathbf{x}_{t-1}$. Denoising diffusion implicit models (DDIM) [DDIM] is a non-Markov diffusion model where each forward process is conditioned on both $\mathbf{x}_{0}$ and $\mathbf{x}_{t-1}$. Specifically, the forward posterior is defined as:
@@ 8.2 Denoising Diffusion Implicit Models
where $\bm{f}_{\theta}(\mathbf{x}_{t},t)=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$. For more detailed proof and derivation of DDIM, see [DDIM].
8.3 Derivation of DDIM Inversion
@@
9 Reverse Scoring Problem
As shown in Tables 9 and 4 (in the main paper), we observe a performance drop when the number of diffusion steps, $S$, increases in conjunction with NLL-based anomaly scoring. To investigate the underlying cause of this phenomenon, we visualize the NLL distribution across different values of $S$ in Figure 4.
Interestingly, the histogram of NLL exhibits counterintuitive results: as $S$ increases by 500, normal images are assigned lower likelihoods, suggesting that learned normal samples occupy lower-density regions than anomalous ones. This observation aligns with findings in flow-based out-of-distribution (OOD) detection literature [DoGMKnow, OODTP], where a similar tendency arises due to the nature of high-dimensional data distributions. This is known as the reverse-scoring issue, which is widely acknowledged in dealing with high-dimensional space.